HUBBLE SPACE TELESCOPE DAILY REPORT #2161 PERIOD COVERED: 0000Z (UTC) 07/08/98 - 0000Z (UTC) 07/09/98 Daily Status Report as of 190/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make snapshot survey observations of the ultraluminous infrared galaxies QDOTULIR06 IR9297-04 and QDOTULIR20. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.4 Completed NIC/1/2/3 7820 (The Metallicity Distribution of the Globular Cluster Systems of Giant Elliptical Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to study the metallicity distribution in the globular clusters contained in the large elliptical galaxy NGC-4874. This is an ideal use of infrared data to significantly augment existing optical data. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed NIC/1 7902 (Photometric Monitoring, Photometric Monitoring-II) The Near Infrared Camera (Camera One) was used to make observations of absolute standard, solar analog star P330E-Cal as part of a photometric monitoring program. Observations were made at several different wavelengths spanning the NICMOS wavelength range in each camera to monitor the photometric stability of the NICMOS cameras as a function of wavelength. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2 7320 (Distribution of Forbidden Iron [Fe-II] Emission at the Galactic Center) The Near Infrared Camera (Cameras One and Two) was used to make observations of the distribution of the forbidden Iron (Fe-II) emissions from the center of our Milky Way galaxy. A conspicuous feature known as the mini-cavity appears as a hole in the distribution of ionized gas within a few arcseconds of the dynamical center of the Galaxy. The observations were executed as scheduled, and no anomalies were reported. 1.7 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two, and Three) were used to make observations of the rich star cluster H11 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.8 Completed STIS/CCD 7730 (Dark Matter at Very Large Radii in M-31) The Space Telescope Imaging Spectrograph (CCD) was used to study Q0038+327 as its light passed through the edge of Galaxy M-31. This is to probe the dark matter distribution of M-31 far from the center by observing intervening absorption from the Mg-II doublet {lambda 2800} in the spectrum of the background QSO located in a geometrically optimal position. A wavelength calibration image was also taken. As described in HSTAR 6574 and in 2.1, a reacquisition for this proposal failed and the take data flag was down for a portion. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.9 Completed Two Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. As described in HSTAR 6574 and in 2.1, a reacquisition for the second iteration of this proposal failed and the take data flag was down for a portion. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.10 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in HSTAR 6574 and in 2.1, a reacquisition for the first iteration of this proposal failed and the take data flag was down for a portion. Otherwise, the observations were completed as planned, and no further problems were reported. 1.11 Completed NIC/1/2 7454 (The Stellar Content of Powerful Radio Galaxies -- A NICMOS Survey of 0.8 < z < 1.8 3CR Sources) The Near Infrared Camera (Cameras One and Two) was used to make survey observations of the powerful radio galaxy 3C266. The study of distant radio galaxies is complicated by their powerful active nucleus, whose emission dominates at rest-frame ultraviolet wavelengths, producing the observed peculiar, elongated morphologies; the underlying stellar component is lost in the glare. As described in HSTAR 6575 and in 2.1, a reacquisition, occurring about three-and-one-half hours into this seven-and-one-half-hour proposal, failed and the take data flag was down for a portion. Otherwise, the observations were completed as planned, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 5 HSTAR 6574 documents a failed reacquisition at 189/194946Z due to the FGS-1 scan steps limit being exceeded. The proposals described in 1.8, 1.9, and 1.10 were affected. HSTAR 6575 describes another failed reacquisition at 190/012948Z, again when the FGS-1 scan steps limit was exceeded. The proposal described in 1.11 was affected as was an iteration of WF/PC 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) which appears during the period of tomorrow's report. 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared once. The DF-224 SPC memory was successfully dumped to CCS strings A and B at 190/0345Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH